OpenStudy (anonymous):

The mean length of six rods is 44.2cm. The mean length of 5 of them is 46cm.How long is the sixth rod?

OpenStudy (anonymous):

let sixth rod length be \(l_6\)\[\frac{l_1+l_2+...+l_5+l_6}{6}=44.2\]\[\frac{l_1+l_2+...+l_5}{5}=46\]

OpenStudy (anonymous):

subistitute the value of \(l_1+l_2+...+l_5\) from second equation in first one to isolate \(l_6\)

Directrix (directrix):

The length of the 6th rod is the total lengths of the 6 rods - the total lengths of the 5 rods which is: 6*44.2 - 5*46 = 35.2 cm for the 6th rod

OpenStudy (anonymous):

but how did u know that it is the total?

OpenStudy (anonymous):

how did u know the sixth rod is the total length of the sixth rod?

OpenStudy (anonymous):

the length of the sixth rod is the total length of all six rods i mean*

Directrix (directrix):

The average is the total of the rods divided by the number of rods. So, the average times the number of rods takes you back to the sum of the lengths.

OpenStudy (anonymous):

could u explain more??

OpenStudy (anonymous):

but theur lengths arent the same ya?

OpenStudy (anonymous):

or is it?

Directrix (directrix):

The average of 3 numbers is 4. What is the sum of the numbers?